# Mapping atmospheric features of the planetary-mass brown dwarf SIMP 0136 with JWST NIRISS

### v1

@@ Abstract
In this paper, we analyze James Webb Space Telescope Near Infrared Imager and Slitless Spectrograph time-series spectroscopy data to characterize the atmosphere of the planetary-mass brown dwarf SIMP J01365662+093347. Principal component analysis reveals that 81% of spectral variations can be described by two components, implying that variability within a single rotational phase is induced by at least three distinct spectral regions. By comparing our data to a grid of Sonora Diamondback atmospheric models, we confirm that the time-averaged spectrum cannot be explained by a single model but require a linear combination of at least three regions. Projecting these models onto the principal component plane shows that the overall variability is highly correlated with changes in temperature, cloud coverage, and possibly effective metallicity. We also extract brightness maps from the lightcurve and establish North-South asymmetry in the atmosphere. A combined multidimensional analysis of spectro-photometric variability links the three spectral regions to three atmospheric layers. Forsterite cloud and water abundance at each level form unique harmonics of atmospheric variability observed in different spectral bands. Atmospheric retrievals on the time-averaged spectrum are consistent with an optically thick iron cloud deck beneath a patchy forsterite cloud layer and with the overall adiabatic curve. We also demonstrate two new analysis methods: a regionally-resolved spectra retrieval that relies on multi-wavelength spherical harmonics maps, and a method to constrain brightness maps using Doppler information present in the spectra. Future observations of variable brown dwarfs of higher spectral resolution or spanning multiple rotations should help break mapping degeneracy.
show]roman.akhmetshyn@mail.mcgill.ca
1 Introduction
@@
Rotational variability is common among brown dwarfs (Artigau et al., 2009; Radigan et al., 2012; Buenzli et al., 2014; Radigan et al., 2014; Metchev et al., 2015; Lew et al., 2016; Artigau, 2018), planetary mass objects (Biller et al., 2015, 2018) and directly imaged companions (Zhou et al., 2016; Naud et al., 2017; Zhou et al., 2020a, b), especially those of young age and low surface gravity (Vos et al., 2018, 2023). Changes in lightcurve morphology are commonly attributed to the dynamic formation and dissipation of clouds, radiative localized hot spots, intense vertical transport due to chemical instability, or the differential rotation of clouds at different latitudes (Bailer-Jones & Mundt, 2001). Specific morphology during a single rotation may be due to cloud features rotating in and out of view. Outstanding questions include whether the atmospheric structure is dominated by banding, spots, or a mixture of both; and measuring the timescales of atmospheric variability to constrain the driving mechanisms.
@@
SIMP J01365662+0933473 (SIMP 0136 hereafter) is a nearby ( pc; Gaia Collaboration et al. 2023) T2.5 brown dwarf discovered in 2006 at the Observatoire du Mont Megantic with the wide-field near-infrared camera CPAPIR as part of SIMP (Sondage Infrarouge de Mouvement Propre) proper-motion survey (Artigau et al., 2006). Further observations revealed a rotation period of hr, a peak-to-peak -band amplitude of 50 mmag, and significant evolution of lightcurve between each observations (Artigau et al., 2009; Croll et al., 2016; Yang et al., 2016; Plummer et al., 2024). These properties make SIMP 0136 the brightest isolated T dwarf and an ideal target to study brown dwarf atmospheric variability. Measurements of constrained its inclination to ∘ (Vos et al., 2017). Further kinematic measurements identified SIMP 0136 to be part of 200 Myr Carina-Near moving group which allowed to estimate its mass of MJ using brown dwarf evolutionary models (Gagné et al., 2017). The same work also estimated that SIMP 0136 could have already depleted half of its deuterium content. Given its mass at the planet-star boundary, SIMP 0136 has been called a free-floating planet. Multiple ground- and space-based observations confirmed its variability across multiple rotations (Artigau et al., 2009; Croll et al., 2016; Yang et al., 2016; Plummer et al., 2024; McCarthy et al., 2025). General circulation models (GCMs) predict that this variability is mainly caused by patchy cloud layers (Tan & Showman, 2021a). It is believed that different cloud decks form and dissipate at different time scales, rotate in and out of the view and hover over each other in different configurations (Apai et al., 2013; Metchev et al., 2015; Apai et al., 2017; Vos et al., 2023).
@@
Section 2 describes the data and its reduction. In Section 3 we show the principal component analysis of our data. In Section 4 we compare NIRISS time-series spectra to atmospheric models. In Section 5 we describe our atmospheric retrieval and compare with the results from previous works. Section 6 describes different mapping approaches, focusing on the results from spherical harmonic mapping. Section 7 explains our multidimensional study of spectral variability and connects lightcurve morphology with meteorological structure. Section 8 explains how spherical harmonics maps can be used to perform extraction and retrieval of regional spectra. Section 9 proposes to constrain brightness maps with Doppler measurements using JWST spectroscopy.
2 Time variable data
@@
The data for this research were obtained with the Near Infrared Imager and Slitless Spectrograph in the Single Object Slitless Spectroscopy mode (NIRISS/SOSS) onboard the James Webb Space Telescope’s (JWST) as a part of cycle 1 Guaranteed Time Observations program #1209 (PI: Artigau). Observations were carried out from UT 05:08:28 to 09:05:07 on July 22, 2023, 37 hours before JWST Cycle 2 program GO #3548 (PI: Vos) that observed SIMP 0136 in NIRSPec/BOTS (McCarthy et al., 2025). We used the SOSS-Inspired SpectroScopic Extraction (SOSSISSE) pipeline for data reduction, which was previously used and described in Lim et al. (2023).
@@
The reduced data consists of 81 1200 spectra from 0.85 to 2.83 m taken over slightly more than one rotational period (Figure 1). The spectrum changes during the object’s rotation, and this variability is different for each wavelength bin associated with certain molecules, suggesting different meteorological structure at different depths.
3 Principal component analysis of time-varying spectra
@@
We use principal component analysis (PCA) to find orthogonal axes that capture the most variance in the spectrum of SIMP 0136 . Since we have 81 spectra, 81 principal components can perfectly replicate all the information in the data. We found that binning data in time does not increase the fraction of variance captured by the first 2–3 of principal components. On the other hand, binning in wavelength greatly increases the fraction of variance explained by the first 2 principal components. Unbinned data have a lot of uncorrelated noise distributed over many eigenvectors, presumably Poisson noise. Averaging it out helps to focus on astrophysical variability. In the rest of this section we use 0.019 m binning, which reduces the number of wavelengths to 93, while the number of principal components remains at 81.
@@
About 81% of the spectral variability can be explained by just two principal components. This justifies projecting our data onto the principal component plane (PCP) made of the first two PCs. According to Cowan et al. (2011), the number of dimensions explaining brightness variability determines the minimum number of distinct spectral regions contributing to the variations: . These spectral regions could represent different cloud decks or clear patches.
@@
The PCP is a useful tool for interpreting the time evolution of spectra (Fuhrmann, 1999; Cowan et al., 2013). In Figure 2, each datum point represents a full spectrum at a given time. From Cowan & Strait (2013), if different spectral regions combine linearly to determine the planet’s overall spectrum, then the pure spectrum of each region must define vertices of a simplex that encloses the data. A simplex is a polyhedron with vertices. For our 2-dimensional PCP a simplex is a triangle. However, as seen in Figure 2, the projected spectra hardly represent a triangular structure and would be better enclosed with a quadrilateral. This suggests that we have 4 end-member spectra.
4 Comparison with Sonora Diamondback Atmospheric Models
@@
In this section we follow a similar approach to Vos et al. (2023) and McCarthy et al. (2024), exploring different possible combinations and number of models that best fit our NIRISS observations. We utilize Sonora family Diamondback spectra models — a new grid of atmospheric models for substellar objects with a range of parameters like temperature, cloud coverage and metallicity (Morley et al., 2024a). We use a limited temperature grid range based on our prior knowledge of SIMP 0136. We also do not constrain gravitational acceleration, because previous studies report values between 100 and 316 cm s-2, see Table 1 for grid parameter values that we explored.
@@
Rescaling the observed spectrum to the top of the atmosphere flux requires knowledge of the radius of the object and the distance to it. While distance is precisely known from Gaia EDR3 (Gaia Collaboration et al., 2023), radius estimations for brown dwarfs are usually inconsistent (Burrows et al., 2011). Atmospheric modeling with BT-Settl (Allard et al., 2013) predict SIMP 0136 to be 0.82 RJup, while evolutionary model SM08 (Saumon & Marley, 2008) given the age estimation provides 1.22 RJup (Gagné et al., 2017). In all our calculations, we adopt the latter, as it gives the most comparable fluxes to Diamondback spectra for the most probable T range. A decrease in radius would greatly increase the emitted flux, thus increasing spectrum-inferred brightness temperature and T of the best-fit Diamondback model above widely accepted values.
@@
First, we search for a single model that best fits the time-averaged spectrum. This corresponds to a model with T K, gravitational acceleration g cm s-2, f, and metallicity of 0.5. To improve the accuracy, we then create finer-grid interpolations around the best model. The best interpolated spectrum has the following parameters: T K, g cm s-2, f, and metallicity of 0.2. As seen in Figure 3, the interpolated model poorly fits the region around 2.0 – 2.5 m, the 1.25 m and 1.65 m peak, and 1.4 m water absorption band. None of the other single grid models described well the time-average spectrum or spectrum at any timestamp. This suggests that the observed spectra are linear combinations of multiple distinct spectra.
@@
We used the Bayesian Information Criterion (BIC) to analyze how many models are needed to best describe our data without over-fitting. The fitting was performed iteratively: starting with the best-fitting single interpolated model, we added additional models and recalculated weights one at a time to improve the overall fit. Our The NIRISS spectrum is incredibly precise, and to adequately apply the BIC, we inflated the uncertainties to match the standard deviation of the spectral variability over time. The number of parameters is , and the BIC is minimized for . This could mean that the observed time-average spectrum is an admixture of spots that have different temperatures and chemistry. An atmosphere explained by 6 spectral models is in agreement with previous PCA results, where majority of the variance is explained by 2 components, and thus there are spectral regions that change in time. Combined PCA and BIC analyses are illustrated in Appendix A. However, without inflating the observational uncertainties, no linear combination of models provides an acceptable match to the data, suggesting that the current models are missing some physical processes and highlighting the need for further refinement.
@@
We constrain the number of admixture components to three based on the results of the PCA and find a best-fit combination for matching the time-averaged spectrum. Physical parameters of these 3 models are: 1000, 1200, and 1300 K; 3, 2, and 8 fsed, 100, 100, and 316 cm s-2; -0.5, +0.5, and +0.5 [M/H]. The integrated Stefan-Boltzmann temperature of the admixture is 1168 K. As shown by the red curve in Figure 3, this reduced admixture model performs only marginally worse than the more complex six-component fit to the time-averaged spectrum. We then use this 3-model combination to fit the time-varying spectra. In Figure 4, we assess whether the observed spectroscopic variability can be captured by the three-model fit. While it slightly underestimates the flux at certain phases, the overall morphology of the variability is well reproduced.
5 Atmospheric retrievals and T-P profiles
@@
We performed atmospheric retrievals on the time-averaged spectrum using the Brewster framework (Burningham et al., 2017, 2021). The temperature profile was parameterized following Madhusudhan & Seager (2009). Our model accounts for the opacity contributions of H2O, CO, CO2, CH4, NH3, CrH, FeH, Na, K, which are spectroscopically active in the near-infrared and have been previously identified as key absorbing species in the spectra of L/T transition objects. Additionally, continuum opacities from collision-induced absorption and Rayleigh scattering due to H2, He, and CH4 are included. Previous retrieval analyses of SIMP 0136 (Vos et al., 2023) indicate the archival 1–15 spectral data are best described by a patchy, high-altitude forsterite/enstatite (Mg2SiO4/MgSiO3) cloud overlaying a deeper, optically thick iron (Fe) cloud; we follow this prescription.
@@
Figure 5 presents the retrieved temperature profile and cloud distributions, compared to Diamondback models (the same used in Figure 3) and phase-equilibrium condensation curves. The retrieved temperature profile (brown curve) shows an adiabatic slope below 1 bar, in good agreement with retrieval results of Vos et al. (2023) that combine 1–2.5 m Infrared Telescope Facility /SpeX Prism spectrum, a 2.5–5 m AKARI/Infrared Camera spectrum, and a 5–15 m Spitzer/IRS spectrum. Self-consistent grid models of the best-fit 3-spectra admixture model are also plotted for reference with Teff = 1000, 1200, 1300 K, g = 100, 100, 316, and = 3, 2, 8. While being broadly consistent, our retrieved profile is more isothermal at high altitudes and slightly cooler at 1–10 bars. This discrepancy may be attributed to the limited flux contribution from the upper atmosphere in the near-infrared bands, resulting in weaker constraints on the temperature structure at these altitudes.
6 Mapping

### v2

@@ Abstract
In this paper, we analyze James Webb Space Telescope Near-Infrared Imager and Slitless Spectrograph time-series spectroscopy data to characterize the atmosphere of the planetary-mass brown dwarf SIMP J01365662+093347. Principal component analysis reveals that 81% of spectral variations can be described by two components, implying that variability within a single rotational phase is induced by at least three distinct spectral regions. By comparing our data to a grid of Sonora Diamondback atmospheric models, we confirm that the time-averaged spectrum cannot be explained by a single model but requires a linear combination of at least three regions. Projecting these models onto the principal component plane shows that the overall variability is highly correlated with changes in temperature, cloud coverage, and possibly effective metallicity. We also extract brightness maps from the lightcurve and establish north–south asymmetry in the atmosphere. A combined multidimensional analysis of spectrophotometric variability links the three spectral regions to three atmospheric layers. Forsterite cloud and water abundance at each level form unique harmonics of atmospheric variability observed in different spectral bands. Atmospheric retrievals on the time-averaged spectrum are consistent with an optically thick iron cloud deck beneath a patchy forsterite cloud layer and with the overall adiabatic curve. We also demonstrate two new analysis methods: a regionally resolved spectra retrieval that relies on multiwavelength spherical harmonics maps, and a method to constrain brightness maps using Doppler information present in the spectra. Future observations of variable brown dwarfs of higher spectral resolution or those spanning multiple rotations should help break the mapping degeneracy.
I Introduction
@@
Rotational variability is common among brown dwarfs (Artigau et al., 2009; Radigan et al., 2012; Buenzli et al., 2014; Radigan et al., 2014; Metchev et al., 2015; Lew et al., 2016; Artigau, 2018), planetary mass objects (Biller et al., 2015, 2018), and directly imaged companions (Zhou et al., 2016; Naud et al., 2017; Zhou et al., 2020a, b), especially those of young age and low surface gravity (Vos et al., 2018, 2023). Changes in lightcurve morphology are commonly attributed to the dynamic formation and dissipation of clouds, radiative localized hot spots, intense vertical transport due to chemical instability, or the differential rotation of clouds at different latitudes (Bailer-Jones & Mundt, 2001). Specific morphology during a single rotation may be due to cloud features rotating in and out of view. Outstanding questions include whether the atmospheric structure is dominated by banding, spots, or a mixture of both, and what the main driving mechanism is that causes these structures to vary in time. One way to constrain that is to measure the timescales of atmospheric variability.
@@
SIMP J01365662+0933473 (SIMP 0136 hereafter) is a nearby ( pc; Gaia Collaboration et al. 2023) T2.5 brown dwarf discovered in 2006 at the Observatoire du Mont-Mégantic with the wide-field near-infrared camera CPAPIR as part of the Sondage Infrarouge de Mouvement Propre (SIMP) proper-motion survey (Artigau et al., 2006). Further observations revealed a rotation period of hr, a peak-to-peak -band amplitude of 50 mmag, and significant evolution of the lightcurve between each observations (Artigau et al., 2009; Croll et al., 2016; Yang et al., 2016; Plummer et al., 2024). These properties make SIMP 0136 the brightest isolated T dwarf and an ideal target to study brown dwarf atmospheric variability. Measurements of constrained its inclination to ∘ (Vos et al., 2017). Further kinematic measurements identified SIMP 0136 to be part of the 200 Myr Carina-Near moving group, which allowed estimation of its mass of MJ using brown dwarf evolutionary models (Gagné et al., 2017). The same work also estimated that SIMP 0136 could have already depleted half of its deuterium content. Given its mass at the planet–star boundary, SIMP 0136 has been called a free-floating planet. Multiple ground- and space-based observations confirmed its variability across multiple rotations (Artigau et al., 2009; Croll et al., 2016; Yang et al., 2016; Plummer et al., 2024; McCarthy et al., 2025). General circulation models (GCMs) predict that this variability is mainly caused by patchy cloud layers (Tan & Showman, 2021a). It is believed that different cloud decks form and dissipate at different timescales, rotate in and out of view, and hover over each other in different configurations (Apai et al., 2013; Metchev et al., 2015; Apai et al., 2017; Vos et al., 2023).
@@
Section II describes the data and its reduction. In Section III, we show the principal component analysis (PCA) of our data. In Section IV, we compare NIRISS time-series spectra to atmospheric models. In Section V, we describe our atmospheric retrieval and compare with the results from previous works. Section VI describes different mapping approaches, focusing on the results from spherical harmonic mapping. Section VII explains our multidimensional study of spectral variability and connects lightcurve morphology with meteorological structure. Section VIII explains how spherical harmonics maps can be used to perform extraction and retrieval of regional spectra. Section IX proposes to constrain brightness maps with Doppler measurements using James Webb Space Telescope (JWST) spectroscopy.
II Time variable data
@@
The data for this research were obtained with the Near-Infrared Imager and Slitless Spectrograph in the Single Object Slitless Spectroscopy (NIRISS/SOSS) mode on board JWST as a part of the Cycle 1 Guaranteed Time Observations program #1209 (PI: Artigau). Observations were carried out from UT 05:08:28 to 09:05:07 on 2023 July 22, 37 hr before JWST Cycle 2 program GO #3548 (PI: Vos) that observed SIMP 0136 in NIRSpec/BOTS (McCarthy et al., 2025). We used the SOSS-Inspired Spectroscopic Extraction (SOSSISSE) pipeline for data reduction, which was previously used and described in Lim et al. (2023).
@@
The reduced data consist of 81 spectra from 0.85 to 2.83 m taken over slightly more than one rotational period (Figure 1). The spectrum changes during the object’s rotation, and this variability is different for each wavelength bin associated with certain molecules, suggesting different meteorological structure at different depths.
III Principal component analysis of time-varying spectra
@@
We use principal component analysis (PCA) to find orthogonal axes that capture the most variance in the spectrum of SIMP 0136. Since we have 81 spectra, 81 principal components can perfectly replicate all the information in the data. We found that binning data in time does not increase the fraction of variance captured by the first two to three principal components. On the other hand, binning in wavelength greatly increases the fraction of variance explained by the first two principal components. Unbinned data have a lot of uncorrelated noise distributed over many eigenvectors, presumably Poisson noise. Averaging it out helps to focus on astrophysical variability. In the rest of this section, we use 0.019 m binning, which reduces the number of wavelengths to 93, while the number of principal components remains at 81.
@@
About 81% of the spectral variability can be explained by just two principal components. This justifies projecting our data onto the principal component plane (PCP) made of the first two principal components. According to Cowan et al. (2011), the number of dimensions explaining brightness variability determines the minimum number of distinct spectral regions contributing to the variations: . These spectral regions could represent different cloud decks or clear patches.
@@
The PCP is a useful tool for interpreting the time evolution of spectra (Fuhrmann, 1999; Cowan et al., 2013). In Figure 2, each datum point represents a full spectrum at a given time. From Cowan & Strait (2013), if different spectral regions combine linearly to determine the planet’s overall spectrum, then the pure spectrum of each region must define vertices of a simplex that encloses the data. A simplex is a polyhedron with vertices. For our two-dimensional PCP, a simplex is a triangle. However, as seen in Figure 2, the projected spectra hardly represent a triangular structure and would be better enclosed with a quadrilateral. This suggests that we have four end-member spectra.
IV Comparison with Sonora Diamondback Atmospheric Models
@@
In this section, we follow a similar approach to Vos et al. (2023) and McCarthy et al. (2024), exploring different possible combinations and numbers of models that best fit our NIRISS observations. We utilize the Sonora family Diamondback spectra models — a new grid of atmospheric models for substellar objects with a range of parameters such as temperature, cloud coverage, and metallicity (Morley et al., 2024a). We use a limited temperature grid range based on our prior knowledge of SIMP 0136. We also do not constrain gravitational acceleration, because previous studies report values between 100 and 316 m s-2; see Table 1 for the grid parameter values we explored.
@@
Rescaling the observed spectrum to the top of the atmosphere flux requires knowledge of the radius of the object and the distance to it. While distance is precisely known from Gaia Early Data Release 3 (Gaia Collaboration et al. 2023), radius estimations for brown dwarfs are usually inconsistent (Burrows et al., 2011). Atmospheric modeling with BT-Settl (Allard et al., 2013) predict SIMP 0136 to be 0.82 RJup, while evolutionary model SM08 (Saumon & Marley, 2008), given the age estimation, provides 1.22 RJup (Gagné et al., 2017). In all our calculations, we adopt the latter, as it gives the most comparable fluxes to Diamondback spectra for the most probable T range. A decrease in radius would greatly increase the emitted flux, thus increasing the spectrum-inferred brightness temperature and T of the best-fit Diamondback model above widely accepted values.
@@
First, we search for a single model that best fits the time-averaged spectrum. This corresponds to a model with T K, gravitational acceleration g m s-2, f, and metallicity of 0.5. To improve the accuracy, we then create finer-grid interpolations around the best model. The best interpolated spectrum has the following parameters: T K, g m s-2, f, and metallicity of 0.2. As seen in Figure 3, the interpolated model poorly fits the region around 2.0–2.5 m, the 1.25 m and 1.65 m peaks, and the 1.4 m water absorption band. None of the other single grid models described well the time-averaged spectrum or spectrum at any time stamp. This suggests that the observed spectra are linear combinations of multiple distinct spectra.
@@
We used the Bayesian Information Criterion (BIC) to analyze how many models are needed to best describe our data without overfitting. The fitting was performed iteratively: starting with the best-fitting single interpolated model, we added additional models and recalculated weights one at a time to improve the overall fit. The NIRISS spectrum is incredibly precise, and to adequately apply the BIC, we inflated the uncertainties to match the standard deviation of the spectral variability over time. The number of parameters is , and the BIC is minimized for . This could mean that the observed time-averaged spectrum is an admixture of spots that have different temperatures and chemistry. An atmosphere explained by six spectral models is in agreement with previous PCA results, where a majority of the variance is explained by two components, and thus there are spectral regions that change in time. Combined PCA and BIC analyses are illustrated in Appendix A. However, without inflating the observational uncertainties, no linear combination of models provides an acceptable match to the data, suggesting that the current models are missing some physical processes and highlighting the need for further refinement.
@@
We constrain the number of admixture components to three based on the results of the PCA and find a best-fit combination for matching the time-averaged spectrum. Physical parameters of these three models are: 1000, 1200, and 1300 K; 3, 2, and 8 fsed; 100, 100, and 316 m s-2; and , , and [M/H]. The integrated Stefan–Boltzmann temperature of the admixture is 1168 K. As shown by the red curve in Figure 3, this reduced admixture model performs only marginally worse than the more complex six-component fit to the time-averaged spectrum. We then use this three-model combination to fit the time-varying spectra. In Figure 4, we assess whether the observed spectroscopic variability can be captured by the three-model fit. While it slightly underestimates the flux at certain phases, the overall morphology of the variability is well reproduced.
V Atmospheric retrievals and T-P profiles
@@
We performed atmospheric retrievals on the time-averaged spectrum using the Brewster framework (Burningham et al., 2017, 2021). The temperature profile was parameterized following Madhusudhan & Seager (2009). Our model accounts for the opacity contributions of H2O, CO, CO2, CH4, NH3, CrH, FeH, Na, and K, which are spectroscopically active in the near-infrared and have been previously identified as key absorbing species in the spectra of L/T transition objects. Additionally, continuum opacities from collision-induced absorption and Rayleigh scattering due to H2, He, and CH4 are included. Previous retrieval analyses of SIMP 0136 (Vos et al., 2023) indicate the archival 1–15 spectral data are best described by a patchy, high-altitude forsterite/enstatite (Mg2SiO4/MgSiO3) cloud overlying a deeper, optically thick iron (Fe) cloud; we follow this prescription.
@@
Figure 5 presents the retrieved temperature profile and cloud distributions, compared to Diamondback models (the same used in Figure 3) and phase-equilibrium condensation curves. The retrieved temperature profile (brown curve) shows an adiabatic slope below 1 bar, in good agreement with retrieval results of Vos et al. (2023) that combine 1–2.5 m Infrared Telescope Facility/SpeX Prism spectrum, a 2.5–5 m AKARI/Infrared Camera spectrum, and a 5–15 m Spitzer/Infrared Spectrograph spectrum. Self-consistent grid models of the best-fit three-spectra admixture model are also plotted for reference with Teff = 1000, 1200, 1300 K, g = 100, 100, 316, and = 3, 2, 8. While being broadly consistent, our retrieved profile is more isothermal at high altitudes and slightly cooler at 1–10 bars. This discrepancy may be attributed to the limited flux contribution from the upper atmosphere in the near-infrared bands, resulting in weaker constraints on the temperature structure at these altitudes.
VI Mapping